OpenStudy (anonymous):

Solve e^(2x) -2e^x + 1 = 0.

OpenStudy (sidsiddhartha):

hint: \[(a-b)^2=a^2-2ab+b^2\]

OpenStudy (sidsiddhartha):

using above identity u can write \[e^{2x}-2e^x+1=(e^x-1)^2\] so \[e^x-1=0\] u can do it from here

OpenStudy (anonymous):

Do you get him?

OpenStudy (anonymous):

I need a full answer please

OpenStudy (sidsiddhartha):

\[e^x-1=0\rightarrow e^x=1\] now take log on both sides \[\log[e^x]=\log1\rightarrow x=0\]